OpenStudy (anonymous):

if m over 3 minus m - 1 over 2 =1,then m =....... help

OpenStudy (bahrom7893):

m/3 - m - 1/2 = 1 multiply everything by 6: 6m/3 - 6m - 6/2 = 6 simplify: 2m - 6m - 3 = 6 add 3 to both sides: 2m - 6m - 3 + 3 = 6 + 3 simplify: -4m = 9 divide both sides by -4: -4m/-4 = 9/-4 simplify: m = -9/4

OpenStudy (anonymous):

i mean m minus 1 divide by 2

OpenStudy (bahrom7893):

m/3 - (m-1)/2 = 1 multiply everything by 6: 6m/3 - 6(m-1)/2 = 6 simplify: 2m - 3(m-1) = 6 open up the parenthesis: 2m - 3m + 3 = 6 subtract 3 from both sides: 2m-3m+3-3=6-3 simplify: -m = 3 multiply both sides by -1: m = -3

OpenStudy (anonymous):

ok thanks sorry for the mistakes

OpenStudy (anonymous):

can i ask somethin ...why the (m-1) become 3m+3

OpenStudy (bahrom7893):

i am multiplying it by -3 So whatever u had, u multiply by 3 and make the sign negative.

OpenStudy (anonymous):

oook